OpenStudy (x__hazel__x):

3. A marble is drawn from a bag containing 3 red marbles and 2 blue marbles and a number cube with faces numbered 1 through 6 is tossed. What is the probability that the marble drawn is blue and the number cube shows a number less than 3? A. 11/15 B. 1/4 C. 1/6 D. 2/15

OpenStudy (fortytherapper):

How many total marbles are there?

OpenStudy (x__hazel__x):

5

OpenStudy (fortytherapper):

Alright, so for marbles, we have a total of 5 and for the dice, it's a total of 6 The question says for a blue marble to be drawn, so we can put \[\frac{ 2 }{ 5 }\] 2 represents the blue marbles and 5 represents the total marbles. Now, how would we do the dice in a fraction?

OpenStudy (x__hazel__x):

hmmm.. so \[\frac{ 4 }{ 6 } ?\]

OpenStudy (fortytherapper):

Almost, they asked for numbers LESS than 3, not 3 or greater

OpenStudy (x__hazel__x):

Ohhhhhhhhhh!!! So, \[\frac{ 2 }{ 6 }\]

OpenStudy (fortytherapper):

Exactly!

OpenStudy (fortytherapper):

Now the last step that we do is multiply (And simplify if needed ofcourse) \[(\frac{ 2 }{ 6 })(\frac{ 2 }{ 5 }) = ?\]

OpenStudy (x__hazel__x):

\[\frac{ 2 }{ 6 } \times \frac{ 2 }{ 5 } = \frac{ 2 }{ 15 }\]

OpenStudy (fortytherapper):

Perfect

OpenStudy (x__hazel__x):

Thank you so much!!

OpenStudy (fortytherapper):

You're welcome =)